# SpatialRGPT: Grounded Spatial Reasoning in Vision-Language Models

### v1

@@ 1 Introduction
Understanding spatial arrangements in both 2D [1, 2] and 3D [3] spaces is crucial for accurately interpreting complex visual environments. Despite the impressive advancements in Vision Language Models (VLMs) across a variety of tasks such as image classification [4], captioning [5], object detection [6], and document parsing [7], etc., these models still face significant challenges with spatial reasoning. This includes difficulties [8, 9] in distinguishing simple spatial concepts like "left" and "right," "above" and "below," as well as more complex relationships such as "behind" and "in front," "inside" and "outside," and "near" and "far." The ability to comprehend and reason about these spatial relationships is fundamental not only for visual understanding, but also for enabling practical applications in fields like robotics [10, 11] and augmented reality [12], where precise spatial awareness is crucial for tasks such as navigation [13], manipulation [10], and interaction with real-world environments [14].
@@
Recently, SpatialVLM [15] has advanced VLMs’ spatial reasoning capabilities by introducing a comprehensive data generation pipeline that enables large-scale training with spatially-aware visual question answering (VQA) tasks. This approach is based on the hypothesis that the limited spatial reasoning capabilities of current VLMs are due to a lack of 3D spatial knowledge in their training data. However, two critical challenges remain. First, effective spatial reasoning requires VLMs to accurately parse regional information, particularly the regions of object instances, whereas most existing VLMs are primarily designed to understand the global context of an image. When an image contains numerous instances, it becomes challenging to prompt the model to reason about the spatial relations between specific instances. This is because most VLMs function as global image parsers and do not support specifying regions for which users want to understand spatial relationships. Second, accurately perceiving spatial relations such as direction and distance cannot rely solely on RGB pixel data. Thus, the architecture needs to incorporate 3D inputs, such as depth information.
@@
In this work, we propose SpatialRGPT, which leverages a data curation pipeline, along with a region and 3D-aware visual encoder architecture to significantly improve the spatial reasoning capability of VLMs. Our data pipeline automatically generates 3D, region-aware annotations from 2D images at scale by constructing a 3D scene graph for each image, where nodes represent object instances and edges denote spatial relationships. This is achieved through three scalable components: (i) open-vocabulary detection and segmentation for instance extraction, (ii) metric depth estimation, and (iii) camera calibration for projecting objects into 3D space. These scene graphs are subsequently transformed into region-aware spatial QA tasks using both template-based and large language model (LLM)-based approaches. This dual approach provides region-based VLMs with the necessary spatial knowledge and advanced reasoning capabilities to interpret complex environments. We use the collected data to train SpatialRGPT. While SpatialRGPT is designed to support region prompts, it effectively avoids the ambiguity issues found in SpatialVLM. In SpatialVLM, multiple similar objects in an image can confuse caption labels. In contrast, our pipeline naturally handles these scenarios without requiring carefully crafted rules or extensive post-processing.
@@ 3 Method
Segmenting objects is the initial stage of building a scene graph. Our models must satisfy two criteria: (i) object descriptions, e.g., class labels, should adhere to an open-world setting for better generalization; (ii) mask proposals need to be highly accurate, ensuring precise contour outlines. This precision is crucial, as even small deviations can lead to significant inaccuracies in the resulting 3D bounding boxes.
To this end, we first employing an open-vocabulary image tagging model [34] to identify all the object classes present in the image. Next, we use Grounding Dino [35], an open-vocabulary 2D detector to determine the corresponding object bounding boxes. Finally, we apply segmentation models [36] to refine these bounding boxes into precise masks. We do not use existing datasets since they either fall short due to vocabulary limitations, or use polygon annotations [37] or compressed masks [38] for segmentation.
@@
Several studies have explored the recovery of metric depth from a single image. The main challenge is to address the scale ambiguity, and one common approach [39, 40] is to use relative depth along with metric heads fine-tuned on specific metric datasets. However, these methods may tend to overfit the depth scale for particular datasets such as KITTI [41], which makes them less robust for images taken in real-world settings. Recently, Metric3Dv2 [42] takes focal length as input and is trained end-to-end to predict metric depth and surface normals. The model is trained on diverse indoor and outdoor scenes, making it less prone to overfitting to the depth distribution of specific datasets. We adopt Metric3Dv2 as our metric depth estimator and found that Metric3Dv2, together with WildCamera’s camera intrinsic, is more robust for images taken in real-world settings and, due to its joint depth-normal optimization training, the recovered metric geometry is improved, particularly around object edges, which is crucial for obtaining accurate 3D bounding boxes.
@@
Camera calibration includes (i) intrinsic matrix estimation to project depth map to 3D point clouds, and (ii) scene canonicalization so that the common scene relations are described in a uniform space. To estimate the camera intrinsic, we use the WildCamera model [43], which estimates 4 degrees of freedom intrinsic parameters (focal point and focal length in two dimensions) and excels in real-world scenarios due to its scale-awareness and ability to account for image cropping. To convert the camera coordinate of the point cloud into a canonicalized geodetic coordinate system for each scene, we leverage PerspectiveFields [44], which provides per-pixel up-vectors and latitude values that can be transformed into camera extrinsics, such as pitch and roll. Using these, we derive a rotation matrix to convert the point cloud from camera coordinates to geodetic coordinates. We note that while SpatialVLM [15] uses surface segmentation (e.g., "floor," "tabletop") to identify a horizontal plane and then uses the normal axis of this plane to align the point cloud to the horizontal plane, however, this approach is limited by the presence of specific classes, e.g., floors or tables, only in the indoor scenes. Also, if there are not enough points for RANSAC, the plane segmentation may fail.
@@ 3.2 Learning Spatial-aware VLMs from 3D Scene Graph
In this section, we discuss converting the constructed 3D scene graph into textual representations for VLM training. One simple approach is through template-based methods via predefined handcrafted instructions. However, this approach limits the diversity of instructions and hinder the model’s reasoning capabilities. Thus, we employ additional complex QAs to enhance the model’s reasoning ability. Our results Fig. 4) show that blending these two types of data can lead to a generalized and complex spatial reasoning model.
@@ 4.2 Standard Visual-language Benchmarks
In this section, we evaluate whether integrating spatial VQA data and depth information affects performance on other VQA tasks. We compared our models with VILA-7B, which is trained on general VQA datasets, and RegionVILA [16], which includes training on both general and region-level datasets. As shown in Table 2, our variants performed similarly to the baselines and slightly better on the VQA-v2 and GQA datasets. These results align with findings from [15], indicating that VLMs generally underperform on spatial reasoning tasks but can improve with specific spatial VQA training without compromising general VQA performance. This highlights the value of our data curation pipeline.
@@ F.1 Instruction Tuning Data
Captioning Image Paragraph Captioning, MSR-VTT, TextCaps
Reasoning CLEVR, NLVR, VisualMRC
Translation Multi30k
VQA ActivityNet-QA, DocVQA, GQA, iVQA, MSRVTT-QA, MSVD-QA, OCR-VQA, ST-VQA, ViQuAE, VQAv2, Visual Dialog
@@ Appendix I Limitations and Future Work
For the most accurate object detection, oriented bounding boxes (OBB) are preferred over axis-aligned bounding boxes (AABB). As illustrated in Figure X, the dimensions obtained from AABBs can differ from those obtained with OBBs. There are two methods to compute an OBB. A simple method involves calculating the OBB using Principal Component Analysis (PCA) of the object’s convex hull, which provides an approximate minimal bounding box. However, this approximation often lacks accuracy due to the incomplete and noisy nature of point clouds captured from a single view. Furthermore, this method still cannot handle extreme cases when objects are partially elevated. The most precise method involves determining the OBB based on the object’s pose, which is currently challenging due to limitations in obtaining accurate object poses. Future improvements could include integrating available pose estimation approaches; however, currently, there are no open-vocabulary solutions for object pose estimation. This remains an area for future research.

### v2

@@ 1 Introduction
Understanding spatial arrangements in both 2D [1, 2] and 3D [3] spaces is crucial for accurately interpreting complex visual environments. Despite the impressive advancements in Vision Language Models (VLMs) across a variety of tasks such as image classification [4], captioning [5], object detection [6], video understanding [7], and document parsing [8], etc., these models still face significant challenges with spatial reasoning. This includes difficulties [9, 10, 11] in distinguishing simple spatial concepts like "left" and "right," "above" and "below," as well as more complex relationships such as "behind" and "in front," "inside" and "outside," and "near" and "far." The ability to comprehend and reason about these spatial relationships is fundamental not only for visual understanding, but also for enabling practical applications in fields like robotics [12, 13] and augmented reality [14], where precise spatial awareness is crucial for tasks such as navigation [15], manipulation [12], and interaction with real-world environments [16].
@@
Recently, several works  [11, 17] has advanced VLMs’ spatial reasoning capabilities by introducing a comprehensive data generation pipeline that enables large-scale training with spatially-aware visual question answering (VQA) tasks. This approach is based on the hypothesis that the limited spatial reasoning capabilities of current VLMs are due to a lack of 3D/2D spatial knowledge in their training data. However, two critical challenges remain. First, effective spatial reasoning requires VLMs to accurately parse regional information, particularly the regions of object instances, whereas most existing VLMs are primarily designed to understand the global context of an image. When an image contains numerous instances, it becomes challenging to prompt the model to reason about the spatial relations between specific instances. This is because most VLMs function as global image parsers and do not support specifying regions for which users want to understand spatial relationships. Second, accurately perceiving spatial relations such as direction and distance cannot rely solely on RGB pixel data. Thus, the architecture needs to incorporate 3D inputs, such as depth information.
In this work, we propose SpatialRGPT, which leverages a data curation pipeline, along with a region and 3D-aware visual encoder architecture to significantly improve the spatial reasoning capability of VLMs.
@@
Our data pipeline automatically generates 3D, region-aware annotations from 2D images at scale by constructing a 3D scene graph for each image, where nodes represent object instances and edges denote spatial relationships. This is achieved through three scalable components: (i) open-vocabulary detection and segmentation for instance extraction, (ii) metric depth estimation, and (iii) camera calibration for projecting objects into 3D space. These scene graphs are subsequently transformed into region-aware spatial QA tasks using both template-based and large language model (LLM)-based approaches. This dual approach provides region-based VLMs with the necessary spatial knowledge and advanced reasoning capabilities to interpret complex environments. We use the collected data to train SpatialRGPT. While SpatialRGPT is designed to support region prompts, it effectively avoids the ambiguity issues found in SpatialVLM. In SpatialVLM, multiple similar objects in an image can confuse caption labels. In contrast, our pipeline naturally handles these scenarios without requiring carefully crafted rules or extensive post-processing.
@@ 3 Method
Segmenting objects is the initial stage of building a scene graph. Our models must satisfy two criteria: (i) object descriptions, e.g., class labels, should adhere to an open-world setting for better generalization; (ii) mask proposals need to be highly accurate, ensuring precise contour outlines. This precision is crucial, as even small deviations can lead to significant inaccuracies in the resulting 3D bounding boxes. To this end, we first employ an open-vocabulary image tagging model [36] to identify all the object classes present in the image. Next, we use GroundingDino [37], an open-vocabulary 2D detector to determine the corresponding object bounding boxes. Finally, we apply segmentation models [38] to refine these bounding boxes into precise masks. We do not use existing dataset annotations since they either fall short due to vocabulary limitations, or use polygon annotations [39] or compressed masks [40] for segmentation.
@@
Several studies have explored the recovery of metric depth from a single image. The main challenge is to address the scale ambiguity, and one common approach [41, 42] is to use relative depth along with metric heads fine-tuned on specific metric datasets. However, these methods may tend to overfit the depth scale for particular datasets such as KITTI [43] or NYU [44], which makes them less robust for in-the-wild images. Recently, Metric3Dv2 [45] takes focal length as input and is trained end-to-end to predict metric depth and surface normals. The model is trained jointly on diverse indoor and outdoor scenes, making it less prone to overfitting to the depth distribution of specific datasets. We adopt Metric3Dv2 as our metric depth estimator and found that Metric3Dv2 together with WildCamera [46]’s camera intrinsic, is robust for images taken in real-world settings. Additionally, thanks to the joint depth-normal optimization training in Metric3Dv2, the recovered geometry is improved particularly around object edges.
@@
Camera calibration includes (i) intrinsic estimation to back-project depth maps to 3D point clouds, and (ii) scene canonicalization to ensure that scene relations are described in a shared space. To estimate the camera intrinsic, we use the WildCamera model [46], which estimates four DoF intrinsic parameters (focal point and focal length in two dimensions). This model excels in real-world scenarios due to its scale-awareness and ability to detect image cropping. To convert the camera coordinates of the point cloud into a canonicalized geodetic coordinate system for each scene, we leverage PerspectiveFields [47], which provides per-pixel up-vectors and latitude values that can be transformed into camera extrinsics, such as pitch and roll. Using these, we derive a rotation matrix to convert the point cloud from camera coordinates to geodetic coordinates. We note that while SpatialVLM [17] uses surface segmentation (e.g., "floor," "tabletop") to identify a horizontal plane and then uses the normal axis of this plane to align the point cloud to the horizontal plane, this approach is limited by the presence of specific classes, such as floors or tables. Additionally, the plane segmentation may fail if there are not enough points for RANSAC.
@@ 3.2 Learning Spatial-aware VLMs from 3D Scene Graph
In this section, we discuss converting the constructed 3D scene graph into textual representations for VLM training. One simple approach is through template-based methods via predefined handcrafted instructions. However, this approach limits the diversity of instructions and hinder the model’s reasoning capabilities. Thus, we employ additional complex QAs to enhance the model’s reasoning ability. Our results in Fig. 4 show that blending these two types of data can lead to a generalized and complex spatial reasoning model.
@@ 4.2 Standard Visual-language Benchmarks
In this section, we evaluate whether integrating spatial VQA data and depth information affects performance on other VQA tasks. We compared our models with VILA-7B, which is trained on general VQA datasets, and RegionVILA [18], which includes training on both general and region-level datasets.
As shown in Table 2, our variants performed similarly to the baselines and slightly better on the VQA-v2 and GQA datasets. These results align with findings from [17], indicating that VLMs generally underperform on spatial reasoning tasks but can improve with specific spatial VQA training without compromising general VQA performance. This highlights the value of our data curation pipeline.
@@ F.1 Instruction Tuning Data
Captioning Image Paragraph Captioning [66], MSR-VTT [67], TextCaps [68]
Reasoning CLEVR [69], NLVR [70], VisualMRC [71]
Translation Multi30k [72]
VQA ActivityNet-QA [73], DocVQA [74], GQA [75], iVQA [76], MSRVTT-QA [67], MSVD-QA [67], OCR-VQA [77], ST-VQA [78], ViQuAE [79], VQAv2 [80], Visual Dialog [81]
@@ Appendix I Limitations and Future Work
For the most accurate object detection, oriented bounding boxes (OBB) are preferred over axis-aligned bounding boxes (AABB). As illustrated in Figure 10, the dimensions obtained from AABBs can differ from those obtained with OBBs. There are two methods to compute an OBB. A simple method involves calculating the OBB using Principal Component Analysis (PCA) of the object’s convex hull, which provides an approximate minimal bounding box. However, this approximation often lacks accuracy due to the incomplete and noisy nature of point clouds captured from a single view. Furthermore, this method still cannot handle extreme cases when objects are partially elevated. The most precise method involves determining the OBB based on the object’s pose, which is currently challenging due to limitations in obtaining accurate object poses. Future improvements could include integrating available pose estimation approaches. However, currently, there are no open-vocabulary solutions for object pose estimation, so this remains an area for future research.